# reference.bib

### v1

@@ 7. Relation to Sugano’s Theorem
with each $f^{(i)}\in\mathbb{C}$, because $H(\xi)_{\infty}\cong SO(n+1)$ is compact. We also note here that $f(u_{i})=f^{(i)}$ for all $i=1,\cdots,h$, as we can see by the way these $f^{(i)}$ are defined in the proof of [Shimura_Euler_Product, Lemma 10.8].
Now, for any two simultaneous eigenfunctions $f_{i},f_{j}$ of the Hecke algebras defined by the pairs $\mathcal{H}_{p}^{\prime}=(H(\xi)_{p},M(g_{f},\xi)_{p})$ for all $p\notin\mathcal{P}$, their inner product is defined via the formula:
@@
where $\Gamma^{k}=H(\xi)_{\mathbb{Q}}\cap u_{i}Du_{i}^{-1}$ and $\nu\left(\Gamma^{k}\right)=\#(\Gamma^{k})^{-1}$, as in [shimura2004arithmetic, (17.23)] (here we again use the fact that $H(\xi)_{\infty}$ is compact).
As $e(\xi)_{i}=\#\{H(\xi)_{\mathbb{Q}}\cap M(u_{i}g_{f};\xi)\}$, we have that $e(\xi)_{i}=\nu\left(\Gamma^{i}\right)^{-1}$, which also gives
@@
It is now possible to choose a basis of orthogonal Hecke eigenforms $\{f_{1},f_{2},\cdots,f_{h}\}$ (we necessarily have $h$ eigenforms) with respect to the above inner product. This is true, because the Hecke algebra defined by $\mathcal{H}_{p}^{\prime}$ is commutative and consists of self-adjoint operators for all $p\notin\mathcal{P}$ (cf. [shimura2004arithmetic, Proposition 17.14]). Hence, we get the expression:
$D_{F,\xi_{i},\mathcal{P}}\left(s+k-\frac{n+2}{2}\right)=\mu(\xi)^{-1}(\textup{zeta})^{-1}L_{\mathcal{P}}(F;s)\sum_{j=1}^{h}\nu\left(\Gamma^{i}\right)\frac{\mu(\xi)}{e(\xi)_{i}^{-1}}f_{j}(u_{i})L_{\mathcal{P}}\left(\overline{f_{j}};s+\frac{1}{2}\right)^{-1}A_{f_{j}},$
which after the simplifications becomes:
$D_{F,\xi_{i},\mathcal{P}}(s)=(\textup{zeta})^{-1}\left(s-k+\frac{n+2}{2}\right)L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}f_{j}(u_{i})L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}A_{f_{j}}.$
We note $A_{f_{j}}$ is $A_{g}$ with $g(u_{i})=\frac{\overline{f_{j}(u_{i})}}{e(\xi)_{i}}\mu(\xi)^{-1}$. Hence, we arrive at the following theorem:
@@
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. Let $\mathcal{P}$ be a finite set of primes, containing the primes described in [sugano, Theorem 1] and all the prime divisors of $D$. Let $F\in S_{k}(\Gamma_{S})$ be a simultaneous eigenfunction for the Hecke algebra $\mathcal{H}_{p}$, defined by the pair $(G_{p},K_{p})$ for all $p\notin\mathcal{P}$ and let $P_{k,D,r}$ denote the Poincaré series of Section 4. Let also $\{u_{i}\}_{i=1}^{h}$ denote a complete system of representatives for $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ such that $u_{i,\infty}=1$ for all $i=1,\cdots,h$ and $\{\xi_{i}\}_{i=1}^{h}$ representatives of $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$ corresponding to $u_{i}$, in the sense described in Proposition 6.2. Let also $\{f_{j}\}_{j=1}^{h}$ denote a complete orthogonal basis of simultaneous eigenfunctions for the pairs $\mathcal{H}_{p}^{\prime}=(H(\xi)_{p},M(g_{f};\xi)_{p})$ for all $p\notin\mathcal{P}$, and denote with $L_{\mathcal{P}}(-,s)$ the standard $L-$function attached to either $F$ or any $f_{j}$, by ignoring the $p-$factors for $p\in\mathcal{P}$. Finally, let $\zeta_{\xi_{i},\mathcal{P}}(s)$ be as in Proposition 6.3. We then have:
@@
$=L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}A_{f_{j}}L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i})\times$
×$\times\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}&\textup{if }n\textup{ even}\end{cases},$
where $A_{g}=\sum_{i=1}^{h}g(u_{i})A(\xi_{i})$, with $g(u_{i})=\frac{\overline{f_{j}(u_{i})}}{e(\xi)_{i}}\mu(\xi)^{-1}$.
@@
By substituting the expression we deduced above in Proposition 6.3, we obtain (here we denote by zeta the function of Theorem 7.1 after $s\longmapsto s-k+(n+2)/2$)
@@
$=(\textup{zeta})^{-1}\times\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}f_{j}(u_{i})L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}A_{f_{j}}=$
$=(\textup{zeta})^{-1}\times L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}A_{f_{j}}L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i}).$
Hence, we would like to explore the connection between $\sum_{i=1}^{h}\zeta_{\xi_{i}}(s-k+n+1)f_{j}(u_{i})$ and $L\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)$.
@@ 8. Explicit Examples
We will now focus our attention to some specific examples of matrices $S$ and corresponding Poincaré series. In particular, we will choose matrices $S$, so that the number of representatives for $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})$ is 11 11 and consider the inner product with the Poincaré series $\mathcal{P}_{k,-q,r}$, with $r\in L$ such that $(-q,r)\in\widetilde{\textup{supp}}(L,\sigma)$. In particular, those choices imply that we can take
@@
as an element of $L_{0}\left[1,\frac{1}{2}\mathbb{Z}\right]$ (because we choose $D=-q$). The main reason we make these choices is the following remark:
Remark 7.
[…]
$\gamma\xi=\begin{pmatrix}*&*&*\\ x&*&y\end{pmatrix}\xi_{0}=\begin{pmatrix}*\\ x+y\end{pmatrix}$
and then $x+y=d>0$. Hence, by the above, $\gamma\in\Gamma(L_{0})\cap G^{*,0}(\mathbb{R})$. Therefore, with the above choices, $\zeta_{\xi,\mathcal{P}}(s)$ of Proposition 6.3 can be written as
@@
8.1. Examples with rank 11 11
We consider the case where $S=2t$ for some $t\geq 1$ with $t$ square-free. This condition is needed so that the lattice $L=\mathbb{Z}$ (and therefore $L_{0}$ and $L_{1}$) is maximal(cf. [eisenstein_thesis, Example 1.6.61.6.6 1.6.61.6.6(ii)]). Now $V_{0}=\mathbb{Q}^{3}$ and the quadratic form of interest is then:
@@ 8.1. Examples with rank 11 11
for $x,y\in V_{0}$, where $S_{0}=\begin{pmatrix}&&1\\ &-2t&\\ 1&&\end{pmatrix}$. Hence, $\phi_{0}$ is represented by $\frac{1}{2}S_{0}$ with respect to the standard basis $e_{1},e_{2},e_{3}$ of $V$.
By [Shimura_Euler_Product, Paragraph 7.3] we have that there exists a quaternion algebra $B$ over $\mathbb{Q}$ such that we can put $V_{0}=B^{\circ}\zeta$, $\phi_{0}[x\zeta]=dxx^{\iota}$, $2\phi_{0}(x\zeta,y\zeta)=dTr_{B/\mathbb{Q}}(xy^{\iota})$, where $B^{\circ}=\{x\in B\mid x^{\iota}=-x\}$ with $\iota$ the main involution of $B$ and $\zeta\in A(V_{0})$ such that $\zeta^{2}=-d$. Here, by $A(V_{0})$ we mean the Clifford algebra of $(V_{0},\phi_{0})$ (see [shimura2004arithmetic, Section 2]). Also, in general, we define $Tr_{B/\mathbb{Q}}(x):=x+x^{\iota}$ for $x\in B$.
Now, [Shimura_Euler_Product, Paragraph 7.3] gives us a way to compute $\zeta$ and $d$. We first need a basis $h_{1},h_{2},h_{3}$ of $V$ such that $\phi_{0}(h_{i},h_{j})=c_{i}\delta_{ij}$ for all $1\leq i,j\leq 3$. In our case, we can make a choice:
@@
We now remind ourselves that $\xi=(1,0,1)^{t}$ and $W=(\mathbb{Q}\xi)^{\perp}$. This then implies that $\phi_{0}[\xi]=1$. From [shimura_diophantine_2006, Paragraph 5.2], we get that $\exists k\in B^{\circ}$ such that $\xi=k\zeta$. Then, if $K:=\mathbb{Q}+\mathbb{Q}k$, we get that $K=\mathbb{Q}(\sqrt{-t})$ because $-t$, which is $-d\phi_{0}[h]$ in the notation there, cannot be a square in $\mathbb{Q}^{\times}$.
We will now use [shimura_diophantine_2006, Theorem 5.7] and the formula $(5.11)$ given there. That, together with the fact that the genus of $L_{0}$ is 11 11 in $V_{0}$ (see Lemma 6.2), will give us the number #$\#\left(L_{0}\left[1,\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})\right)$ by $(5.15\textup{a})$ in [shimura_diophantine_2006, Theorem 5.10]. This will equal $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]$ and we are interested in the cases where this number is 11 11.
@@
We start by first computing the quantities appearing in the statement of [shimura_diophantine_2006, Theorem 5.7]. We have that $\mathfrak{e}$ is the product of all the prime ideals in $\mathbb{Q}$ for which $B$ ramifies. However, for each prime $p$, $\phi_{0}$, viewed as a bilinear form over $\mathbb{Q}_{p}$ is isotropic (the Witt index has to be 11 11 for all primes $p$). Therefore by [Shimura_Euler_Product, Paragraph 7.3], $B$ over $\mathbb{Q}_{p}$ is not a division algebra, hence is isomorphic to $M_{2}(\mathbb{Q}_{p})$, i.e. $B$ splits or is unramified over $p$. Hence, $\mathfrak{e}=\mathbb{Z}$.
Now, $\mathfrak{a}=t\mathbb{Z}$ because $t$ is square-free and for the different $\mathfrak{d}_{K/F}$ we have
$\mathfrak{d}_{K/F}=\begin{cases}2\sqrt{-t}\mathcal{O}_{K}&\textup{ if }-t\nequiv 1\pmod{4}\\ \sqrt{-t}\mathcal{O}_{K}&\textup{ if }-t\equiv 1\pmod{4}.\end{cases}$
We now want to investigate the condition $\mathfrak{c}_{p}^{2}N_{K/F}(\mathfrak{d}_{K/F})_{p}=\mathfrak{a}_{p}\phi_{0}[\xi]\phi_{0}(\xi,L_{0})_{p}^{-2}$ for a prime $p$ for the conductor $\mathfrak{c}$. But, by the above:
$N_{K/F}(\mathfrak{d}_{K/F})=\begin{cases}4t\mathbb{Z}&\textup{ if }-t\nequiv 1\pmod{4}\\ t\mathbb{Z}&\textup{ if }-t\equiv 1\pmod{4},\end{cases}$
and $\phi_{0}[\xi]=1$, $\phi_{0}(\xi,L_{0})=\frac{1}{2}\mathbb{Z}$ and so we get
$\mathfrak{c}_{p}^{2}\cdot\begin{cases}4t\mathbb{Z}_{p}&\textup{ if }-t\nequiv 1\pmod{4}\\ t\mathbb{Z}_{p}&\textup{ if }-t\equiv 1\pmod{4}\end{cases}=4t\mathbb{Z}_{p}.$
Therefore, if $p\neq 2$, $\mathfrak{c}_{p}=\mathbb{Z}_{p}$ and if $p=2$, $\mathfrak{c}_{2}=\mathbb{Z}_{2}$ if $-t\nequiv 1\pmod{4}$ and $\mathfrak{c}_{2}=2\mathbb{Z}_{2}$ otherwise.
@@
Also, $N_{K/F}(\mathcal{O}_{K}^{\times})=\{1\}$ because $K$ is imaginary quadratic and so [×$\left[\mathbb{Z}^{\times}:N_{K/F}(\mathcal{O}_{K}^{\times})\right]=2$. Finally, $\infty$ ramifies and so $\nu=1$. We will now consider specific cases in order to determine the ideal $\mathfrak{a}^{*}$ and the sets $U,U^{\prime}$.
  • $t=1$. Then, $\mathfrak{a}=\mathfrak{e}=\mathfrak{a}^{*}=\mathbb{Z}$ and from the above $\mathfrak{c}=\mathbb{Z}$ as well. Also, $U=\mathcal{O}_{K}^{\times}=\{\pm 1,\pm i\}$ and now
@@
    Clearly ±$\pm 1\in U^{\prime}$ but ±$\pm i\notin U^{\prime}$ because $2=(1+i)^{2}=(1-i)^{2}$. So, $[U:U^{\prime}]=2$. Finally, $\mu=0$ because $\mathfrak{a}^{*}\mathfrak{e}=\mathbb{Z}$ and therefore, we get
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=(1/1)\cdot 2^{1-0-1}\cdot 2\cdot 2^{-1}=1,$
    as claimed.
  • $t=2$. Then $\mathfrak{a}=2\mathbb{Z}$ and we can compute $\mathfrak{a}^{*}=2\mathbb{Z}$. This is true because the only prime factor of $\mathfrak{a}$ is 22 22 and $r_{2}(\xi)=-2\in\frac{1}{2}\mathbb{Z}_{p}$ and $r_{2}(\xi)\notin\mathfrak{C}_{2}$ (if $-2=u^{2}+4w$ with $u,v\in\mathbb{Z}_{2}$, then $2\mid u$ which then implies $2\mid 1$, contradiciton). The definitions of $r_{2}(\xi)$ and $\mathfrak{C}$ can be found in [shimura_diophantine_2006, Paragraph 5.5]. Also, by the above, $\mathfrak{c}=\mathbb{Z}$. Now, $U=\mathcal{O}_{K}^{\times}=\{\pm 1\}$ and
    $U^{\prime}=\{x\in\mathcal{O}_{K}^{\times}\mid x-1\in 2\sqrt{-2}\mathcal{O}_{K,p}\textup{ }\forall p\neq 2\}.$
    But both ±$\pm 1$ belong to $U^{\prime}$ as $\sqrt{-2}$ is a unit in $\mathcal{O}_{K,p}$ for all $p\neq 2$. Hence, $[U:U^{\prime}]=1$. Finally, $\mu=1$ in this case because 2=($2=(\sqrt{-2})^{2}$ is ramified. Therefore:
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=(1/1)\cdot 2^{1-1-1}\cdot 2\cdot 1^{-1}=1.$
  • $t=3$. Here $\mathfrak{a}=3\mathbb{Z}$ and $\mathfrak{a}^{*}=3\mathbb{Z}$ because $r_{3}(\xi)=-\frac{4}{3}\in\frac{1}{3}\mathbb{Z}_{3}$ but $r_{3}(\xi)\notin\mathfrak{C}_{3}$ as $\mathfrak{C}_{3}=\mathbb{Z}_{3}$. By what we said above for the conductor, we get $\mathfrak{c}_{2}=2\mathbb{Z}_{2}$ and $\mathfrak{c}_{p}=\mathbb{Z}_{p}$ for $p\neq 2$. Hence, $\mathfrak{c}=2\mathbb{Z}$. In this case, $\mathcal{O}_{K}^{\times}=\{\pm 1,\pm\omega,\pm\omega^{2}\}$, where $\omega=\frac{1}{2}(1+\sqrt{-3})$. Then,

### v2

@@ 8. Relation to Sugano’s Theorem
with each $f^{(i)}\in\mathbb{C}$, because $H(\xi)_{\infty}\cong\hbox{SO}(n+1)$ is compact. We also note here that $f(u_{i})=f^{(i)}$ for all $i=1,\cdots,h$, as we can see by the way these $f^{(i)}$ are defined in the proof of [22, Lemma 10.8].
Now, for any two simultaneous eigenfunctions $f_{i},f_{j}$ of the Hecke algebras defined by the pairs $\mathcal{H}_{p}^{\prime}=(H(\xi)_{p},M(g_{f},\xi)_{p})$ for all $p\notin\mathcal{P}$, their inner product is defined via the formula
@@
where $\Gamma^{k}=H(\xi)_{\mathbb{Q}}\cap u_{k}Du_{k}^{-1}$ and $\nu\left(\Gamma^{k}\right)=\#\left(\Gamma^{k}\right)^{-1}$, as in [24, (17.23)] (here we again use the fact that $H(\xi)_{\infty}$ is compact).
As $e(\xi)_{i}=\#\{H(\xi)_{\mathbb{Q}}\cap M(u_{i}g_{f};\xi)_{f}\}$, we have that $e(\xi)_{i}=\nu\left(\Gamma^{i}\right)^{-1}$, which also gives
@@
It is now possible to choose a basis of orthonormal Hecke eigenforms $\{f_{1},f_{2},\cdots,f_{h}\}$ for $V(g_{f};\xi)$ with respect to the above inner product (i.e. $\langle f_{i},f_{j}\rangle=\delta_{ij}$ for all $i,j$). This is true because the Hecke algebra defined by $\mathcal{H}_{p}^{\prime}$ is commutative and consists of self-adjoint operators for all $p\notin\mathcal{P}$ (see proof of [24, Proposition 17.14]). Also, by [24, Lemma 17.6, (1)], there is an isomorphism between $V(g_{f};\xi)$ and $\mathbb{C}^{h}$. Therefore, this basis must consist of $h$ eigenforms. Hence, we get the expression
$D_{F,\xi_{i},\mathcal{P}}\left(s+k-(n+2)/2\right)=\mu(\xi)^{-1}(\textup{zeta})^{-1}(s)L_{\mathcal{P}}(F;s)\sum_{j=1}^{h}\nu\left(\Gamma^{i}\right)\frac{\mu(\xi)}{e(\xi)_{i}^{-1}}f_{j}(u_{i})\times\\ \times L_{\mathcal{P}}\left(\overline{f_{j}};s+1/2\right)^{-1}A_{f_{j}},$
which after the simplifications becomes
$D_{F,\xi_{i},\mathcal{P}}(s)=(\textup{zeta})^{-1}\left(s-k+(n+2)/2\right)L_{\mathcal{P}}\left(F;s-k+(n+2)/2\right)\times\\ \times\sum_{j=1}^{h}f_{j}(u_{i})L_{\mathcal{P}}\left(\overline{f_{j}};s-k+(n+3)/2\right)^{-1}A_{f_{j}}.$ (16)
Hence, we arrive at the following Theorem.
@@
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. Let $\mathcal{P}$ be a finite set of primes, containing the primes described in the proof of Theorem 8.1 and all the prime divisors of $D$. Let $F\in S_{k}(\Gamma_{S})$ corresponding to $\bm{F}\in\mathfrak{S}(K_{f})$ with $A(\xi)\neq 0$. Assume $\bm{F}$ is a simultaneous eigenfunction for the Hecke algebra $\mathcal{H}_{p}$, defined by the pair $(G_{p},K_{p})$ for all $p$ and let $\mathcal{P}_{k,D,r}$ denote the Poincaré series of (8). Let also $\{f_{j}\}_{j=1}^{h}\in V(g_{f};\xi)$ denote an orthonormal basis of simultaneous eigenfunctions for the pairs $\mathcal{H}_{p}^{\prime}=(H(\xi)_{p},M(g_{f};\xi)_{p})$ for all $p\notin\mathcal{P}$, $A_{f_{j}}$ as in (15), and denote with $L_{\mathcal{P}}(-,s)$ the standard $L$-function attached to either $F$ or any $f_{j}$, by ignoring the $p$-factors for $p\in\mathcal{P}$. We then have
@@
$L_{\mathcal{P}}\left(F;s-k+(n+2)/2\right)\sum_{j=1}^{h}A_{f_{j}}L_{\mathcal{P}}\left(\overline{f_{j}};s-k+(n+3)/2\right)^{-1}\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i})\times$
$\times\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}&\textup{if }n\textup{ even}\end{cases},$
where $\zeta_{\xi_{i},\mathcal{P}}(s)$ are as in Proposition 7.3.
@@
By substituting the expression we deduced in (16) into (13), we obtain (here we denote by "zeta" the function of Theorem 8.1 after $s\longmapsto s-k+(n+2)/2$)
@@
$(\textup{zeta})^{-1}\times\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}f_{j}(u_{i})L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}A_{f_{j}}$
$(\textup{zeta})^{-1}\times L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\sum_{j=1}^{h}A_{f_{j}}L_{\mathcal{P}}\left(\overline{f_{j}};s-k+\frac{n+3}{2}\right)^{-1}\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i}).\qed$
Hence, we would like to explore the connection between $\sum_{i=1}^{h}\zeta_{\xi_{i}}(s-k+n+1)f_{j}(u_{i})$ and $L\left(\overline{f_{j}};s-k+(n+3)/2\right)$.
@@ 9. Explicit Examples
We will now focus our attention to some specific examples of matrices $S$ and corresponding Poincaré series. In particular, we set $D=-q$ and choose matrices $S$, so that the number of representatives for $L_{0}\left[1,\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})$ is $1$. We therefore take $G$ to be the Poincaré series $\mathcal{P}_{k,-q,r}$, with $r\in L$ such that $(-q,r)\in\widetilde{\textup{supp}}(L,\sigma)$. In particular, those choices imply that we can take
@@
as an element of $L_{0}\left[1,\frac{1}{2}\mathbb{Z}\right]$. Therefore, $\zeta_{\xi,\mathcal{P}}(s)$ of Proposition 7.3 can be written as
@@
9.1. Examples with rank $1$
We consider the case where $S=2t$ for some $t\geq 1$ with $t$ square-free. This condition is needed so that the lattice $L=\mathbb{Z}$ (and therefore $L_{0}$ and $L_{1}$) is maximal (see [21, Example 1.6.6(ii)]). Now $V_{0}=\mathbb{Q}^{3}$ and the quadratic form of interest is then
@@ 9.1. Examples with rank $1$
for $x,y\in V_{0}$, where $S_{0}=\begin{pmatrix}&&1\\ &-2t&\\ 1&&\end{pmatrix}$. Hence, $\phi_{0}$ is represented by $S_{0}/2$ with respect to the standard basis $e_{1},e_{2},e_{3}$ of $V$.
By [24, Paragraph 7.3] we have that there exists a quaternion algebra $B$ over $\mathbb{Q}$ such that we can put $V_{0}=B^{\circ}\zeta$, $\phi_{0}[x\zeta]=dxx^{\iota}$, $2\phi_{0}(x\zeta,y\zeta)=d\textup{Tr}_{B/\mathbb{Q}}(xy^{\iota})$, where $B^{\circ}=\{x\in B\mid x^{\iota}=-x\}$ with $\iota$ the main involution of $B$ and $\zeta\in A(V_{0})$ such that $\zeta^{2}=-d$. Here, by $A(V_{0})$ we mean the Clifford algebra of $(V_{0},\phi_{0})$ (see [24, Chapter 2]). Also, in general, we define $\textup{Tr}_{B/\mathbb{Q}}(x):=x+x^{\iota}$ for $x\in B$.
Now, from [24, Paragraph 7.3], we have a way to compute $\zeta$ and $d$. We first need a basis $h_{1},h_{2},h_{3}$ of $V$ such that $\phi_{0}(h_{i},h_{j})=c_{i}\delta_{ij}$ for all $1\leq i,j\leq 3$. In our case, we can make a choice
@@
We now remind ourselves that $\xi=(1,0,1)^{t}$ and $W=(\mathbb{Q}\xi)^{\perp}$. This then implies that $\phi_{0}[\xi]=1$. From [26, Paragraph 5.2], we get that $\exists k\in B^{\circ}$ such that $\xi=k\zeta$. Then, if $K:=\mathbb{Q}+\mathbb{Q}k$, we get that $K=\mathbb{Q}(\sqrt{-t})$ because $-t$, which is $-d\phi_{0}[\xi]$, cannot be a square in $\mathbb{Q}^{\times}$.
Using now [26, Theorem 5.7] and the formula $(5.11)$ given there, tailored to our situation, we have the following Theorem.
Theorem 9.1.
We define the following quantities:
  • Let $K:=\mathbb{Q}(\sqrt{-t})$ and $c_{K}$ denote the class number of $K$.
  • $\mathfrak{c}$ denotes the ideal of $\mathbb{Z}$, determined by the local conditions
    $\mathfrak{c}_{p}^{2}N_{K/\mathbb{Q}}(\mathfrak{d}_{K/\mathbb{Q}})_{p}=\mathfrak{a}_{p}\phi_{0}[\xi]\phi_{0}(\xi,L_{0})_{p}^{-2},$ (17)
    for all primes $p$, where $\mathfrak{a}=t\mathbb{Z}$ and $\mathfrak{d}_{K/\mathbb{Q}}$ is the different ideal.
[…]
  • Let $p$ be a rational prime. Pick $\epsilon_{p}\in-\det(\phi_{0,p})(\mathbb{Q}_{p}^{\times})^{2}$, which is either a unit or a prime element of $\mathbb{Q}_{p}$ and choose an element $\beta_{p}\in\mathbb{Q}_{p}$ such that $\phi_{0}(\xi,L_{0,p})=\beta_{p}\mathbb{Z}_{p}$. Define then $r_{p}(\xi):=\epsilon_{p}^{-1}\phi_{0}[\xi]\beta_{p}^{-2}$. Define also
    $\mathfrak{C}_{p}:=\{u^{2}+4w\mid u,w\in\mathbb{Z}_{p}\}.$
  • $\mathfrak{a}^{*}$ is the product of the prime factors $p$ of $t$ such that $r_{p}(\xi)\in p^{-1}\mathbb{Z}_{p}$ and $r_{p}(\xi)\notin\mathfrak{C}_{p}$.
  • $\mu$ is the number of prime ideals dividing $\mathfrak{a}^{*}$ and ramified in $K$.
  • $U:=\mathcal{O}_{K}^{\times}$ and $U^{\prime}:=\{x\in\mathcal{O}_{K}^{\times}\mid x-1\in\mathfrak{c}_{p}(\mathfrak{d}_{K/\mathbb{Q}})_{p}\ \forall p\nmid\mathfrak{a}^{*}\}$.
We then have
$H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=c_{K}2^{1-\mu}[U:U^{\prime}]^{-1}N(\mathfrak{c})\prod_{p\mid\mathfrak{c}}\left\{1-\frac{1}{p}[K/\mathbb{Q},\ p]\right\},$ (18)
Proof.
This follows from [26, Theorem 5.7]. In our case, we have, in the notation of the Theorem:
  • The base field $F$ is $\mathbb{Q}$, which has class number $1$.
@@
  • The product of all the prime ideals in $\mathbb{Q}$ for which $B$ ramifies, $\mathfrak{e}$, equals $\mathbb{Z}$. This is true because for each prime $p$, $\phi_{0}$, viewed as a bilinear form over $\mathbb{Q}_{p}$, is isotropic (the Witt index is $1$ for all primes $p$). Therefore by [24, Paragraph 7.3], $B$ over $\mathbb{Q}_{p}$ is not a division algebra, hence is isomorphic to $M_{2}(\mathbb{Q}_{p})$, i.e. $B$ splits (or is unramified) over $p$. Hence, $\mathfrak{e}=\mathbb{Z}$.
  • Because $t$ is square-free, $\mathfrak{a}=t\mathbb{Z}$.
@@
  • Because $K$ is imaginary quadratic, $\infty$ ramifies, so $\nu=1$ and $N_{K/\mathbb{Q}}(\mathcal{O}_{K}^{\times})=\{1\}$, so $\left[\mathbb{Z}^{\times}:N_{K/\mathbb{Q}}(\mathcal{O}_{K}^{\times})\right]=2$.∎
The above Theorem gives us the number $\displaystyle{\#\left(L_{0}\left[1,\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})\right)}$ from Proposition 7.2 and the fact that in this case, $H(\xi)$ is commutative (see proof of [26, Theorem 5.10]). We are interested in the cases when this number is $1$.
For the different $\mathfrak{d}_{K/\mathbb{Q}}$, we have
$\mathfrak{d}_{K/\mathbb{Q}}=\begin{cases}2\sqrt{-t}\mathcal{O}_{K}&\textup{ if }-t\not\equiv 1\pmod{4}\\ \sqrt{-t}\mathcal{O}_{K}&\textup{ if }-t\equiv 1\pmod{4}\end{cases}$.
We now want to determine the ideal $\mathfrak{c}$. But, by the above
$N_{K/\mathbb{Q}}(\mathfrak{d}_{K/\mathbb{Q}})=\begin{cases}4t\mathbb{Z}&\textup{ if }-t\not\equiv 1\pmod{4}\\ t\mathbb{Z}&\textup{ if }-t\equiv 1\pmod{4}\end{cases}$,
and $\phi_{0}[\xi]=1$, $\displaystyle{\phi_{0}(\xi,L_{0})=\frac{1}{2}\mathbb{Z}}$. So, from (17), we get the equation
$\mathfrak{c}_{p}^{2}\cdot\begin{cases}4t\mathbb{Z}_{p}&\textup{ if }-t\not\equiv 1\pmod{4}\\ t\mathbb{Z}_{p}&\textup{ if }-t\equiv 1\pmod{4}\end{cases}=4t\mathbb{Z}_{p}.$
Therefore,
$\mathfrak{c}=\begin{cases}\mathbb{Z}&\textup{ if }-t\not\equiv 1\pmod{4}\\ 2\mathbb{Z}&\textup{ if }-t\equiv 1\pmod{4}\end{cases}$. (19)
We will now consider specific cases in order to determine when the index in (18) is 1.
  • $t=1$. Then, $\mathfrak{a}=\mathfrak{e}=\mathfrak{a}^{*}=\mathbb{Z}$ and from (19) $\mathfrak{c}=\mathbb{Z}$. Now
@@
    Clearly $\pm 1\in U^{\prime}$ but $\pm i\notin U^{\prime}$ because $2\mathcal{O}_{K}=(1+i)^{2}\mathcal{O}_{K}=(1-i)^{2}\mathcal{O}_{K}$. So, $[U:U^{\prime}]=2$. Finally, $\mu=0$ because $\mathfrak{a}^{*}\mathfrak{e}=\mathbb{Z}$ and therefore, we get
    $H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=1\cdot 2^{1-0}\cdot 2^{-1}=1.$
  • $t=3$. Here $\mathfrak{a}=3\mathbb{Z}$ and $\mathfrak{a}^{*}=3\mathbb{Z}$ because $r_{3}(\xi)=-\frac{4}{3}\in\frac{1}{3}\mathbb{Z}_{3}$ but $r_{3}(\xi)\notin\mathfrak{C}_{3}$, as $\mathfrak{C}_{3}=\mathbb{Z}_{3}$. By (19), we get $\mathfrak{c}=2\mathbb{Z}$. In this case, $\mathcal{O}_{K}^{\times}=\{\pm 1,\pm\omega,\pm\omega^{2}\}$, where $\omega=\frac{1}{2}(1+\sqrt{-3})$. Then,